# Cosmology of self-replicating universes in black holes formed by dark matter-seeded stellar collapse

### v1

@@ II Summary
    Here we will propose that most of our universe’s CNS-associated black hole production may actually occur through black hole converting old helium white dwarfs (WDs) into about $0.1~{}{M_{\odot}}$ black holes. This can transmute nearly all baryons in galaxies into black holes over 100 trillion year timescales.
@@
Figure 1: The total entropy production in our universe with associated uncertainties (solid blue lines) as a function of time, using the historic star formation rate and uncertainties reported in Ref. Hopkins and Beacom (2006), and extrapolated to future epochs using future star formation Adams and Laughlin (1997) as described in Sec. IV.2. The dashed black curve shows the amount of galactic stellar material that would be turned into black holes for dark matter with the mass and per-nucleon cross section indicated. This result is shown both in terms of the number fraction (labelled “##\##”) and mass fraction (labelled “mass”) of stars converted. In the absence of dark matter conversion (“no DM conv.”), baryons predominantly form black holes through core collapse supernovae in a roughly constant proportion Sicilia et al. (2022). The calculations in this work account for dark matter conversion to black holes of $0.1-0.5~{}M_{\odot}$ He white dwarfs, which are the predominant site of baryons in galaxies after star formation ceases. Dark matter-induced collapse will also apply to higher mass white dwarfs, neutron stars Gould et al. (1990); Kouvaris (2008); Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010); McDermott et al. (2012); Kouvaris and Tinyakov (2011); Kouvaris (2012); Güver et al. (2014); Bertoni et al. (2013); Bramante et al. (2013); Bell et al. (2013); Bramante and Linden (2014); Bramante (2015); Bramante and Linden (2016); Bramante et al. (2018); Garani et al. (2019); Kouvaris et al. (2018); Kopp et al. (2018); Acevedo and Bramante (2019); East and Lehner (2019); Tsai et al. (2021); Dasgupta et al. (2020); Garani et al. (2021, 2022); Giffin et al. (2022); Ray (2023), planets Acevedo et al. (2021), and other celestial objects, a sub-dominant contribution not included here.
@@ III.1 Cosmological natural selection
Cosmological natural selection is based on the notion that typical universes are those that reproduce best. As detailed in Ref. Smolin (1992), the reproduction of universes is hypothesized to occur via black hole formation, with the assumption that a causally disconnected universe is born beyond the horizon from a bouncing singularity. In each new universe, physical constants slightly vary, presumably due to a dense discretuum of vacua whose minima vary by a small amount due to the bounce dynamics (such as might be seen in something like a string theory landscape). This results in a sum of universes whose greater part is optimized to create black holes. Ref. Smolin (1992) points out, for instance, that if the neutron-proton mass difference were negative, radiative cooling of neutron (as opposed to hydrogen) gas clouds would be too slow and suppress star formation; if it were higher than the observed 1.29 MeV, nucleosynthesis of elements heavier than hydrogen would be slower, which in turn would subdue the formation of later-generation stars from recycled supernova-ejected stellar material. If it were too high, stable deuterium, and thence heavier nuclides, may never be formed in Big Bang nucleosynthesis. Analogous to the CNS mechanism is Darwinian natural selection, with the slight changes in genetic code paralleling the slight changes in physical constants.
@@ III.4 Linking entropy and black hole production
We now propose the following update to CNS in Smolin’s original form, which for clarity here we will call CNS1. In CNS1, about 1% of baryonic mass ends up inside black holes (through conventional core-collapse of superheavy stars); in the new version, more than 75% of the galactic baryonic mass becomes black holes. [^1] As discussed later, this number is likely higher once black hole conversion of $>0.5$ solar mass stars and NSs are accounted for. In CNS1 the criteria by which the framework is evaluated is tied to known black hole production mechanisms, but does not tie into the dynamic timescale and production of entropy in our universe; in the CNS proposed here, timescales for entropy production and creation of new universes come out to be about the same, and this is quantified using stelliferous entropy. Finally, in our proposed reframing of CNS, DM plays a central and intricate role in converting stars into black holes.
@@ IV Conversion of white dwarfs to black holes via dark matter
In this section, we estimate the timescale over which low-mass helium WDs, end products of low-mass main sequence stars, would be converted into black holes by heavy asymmetric dark matter. We assume that dark matter is a Dirac fermion, although as we will see shortly our treatment would apply to bosonic dark matter states as well. For further aspects of dark matter capture in compact stars see the review in Ref. Bramante and Raj (2024).
@@ IV.1 Dark matter-white dwarf dynamics
where this expression sums over rates for dark matter to be captured after $n$ scatters, given by
@@
IV.2 Entropy production and BH formation
To estimate the integrated entropy production rate shown in Figure 2 at time $t$ we generally follow Ref. Bousso et al. (2007), with some modifications. We use the Kroupa IMF Kroupa (2002) given in Sec. III.3, $\xi_{\rm IMF}(M_{\star})$, which weights the entropy production rate per star, $d^{2}s/dN_{\star}dt=L_{\star}/T_{\rm eff}$, where the main sequence star luminosity $L_{\star}\propto M_{\star}^{3.5}$ and the temperature of reprocessed starlight $T_{\rm eff}=20$ meV. This is then convolved with the redshift-dependent star formation rate, $\dot{\rho}_{\star}$, and integrated over the birth times $t^{\prime}$ of stars born prior to $t$ and volume $V_{c}$:
@@ V.1 Alternative CNS measures and dynamics
This brings us to the question of whether non-stellar sources of entropy increase might be considered. In our work we have used the entropy sourced by reprocessed starlight as the predominant non-horizon entropy source. But far more entropy than this is sourced by horizons, $i.e.$ of both BHs and the cosmic horizon of the observable universe. As in Ref. Bousso et al. (2007) we have not considered this horizon entropy in the main text, as its relevance to observers and the birth of universes is unclear. If horizon entropy is considered for some reason, our universe as we know it is enough to convert most of the mass of entropy sources to new universes. This is because in $10^{33}$ yr most baryonic mass would have gone into making galactic mass BHs through dynamical accumulation onto the supermassive black holes (SMBHs) at the centers of galaxies Adams and Laughlin (1997). These eventual SMBHs would be the major sources of horizon entropy. In that scenario, star formation plays a key role, but the role of dark matter is unclear. But if it is part of the equation, models of dark matter seeding SMBHs Feng et al. (2021) become significant. The BH count may also be maximized in these scenarios. In addition, one could have super-massive PBHs that grow faster than regular SMBHs and contributing to the majority of entropy, although these could only make up a small fraction of dark matter owing to constraints from their accretion Green and Kavanagh (2021). Also, these dark matter-as-black holes versions of CNS do not explain star formation.
@@ V.2 Perspective on reccurrence
To wrap up, we discuss measure problems and their relation to CNS. As a prior example of measure problems, let us consider Weinberg’s work on understanding the cosmological constant. In formulating his anthropic bound on the CC, Weinberg had assumed that the fraction of baryons that ends up in galaxies is another parameter (in addition to the CC) that scans on some landscape of vacua, as the anthropically conditioned probability distribution is proportional to the number of observers. But he had raised the possibility that the number of baryons may itself scan Weinberg (2005), in which case one cannot know that our universe is typical. Analogously, one parameter that could scan in the CNS framework is the fraction of baryonic mass that converts to BHs. As we have noted at the outset of our discussion, scanning parameters in this way does not necessarily lead to increased universe fitness.
However, in the context of cosmologies whose future looks similar to their past, let us note that CNS appears to generically circumvent the Boltzmann brain problem. Over timescales of $\gg 10^{10^{66}}$ yr, thermodynamic fluctuations could create low-entropy states that mimic conscious observers with false memories of a briefer evolution of the universe Carroll (2017). The problem is that these “Boltzmann brains” would far outnumber “regular brains” formed through better known evolutionary pathways, implying we must ask whether we are the former. The solution would be to construct a cosmology where Boltzmann brains are atypical. In CNS as we have outlined, the formation of new universes occurs over $10^{14}$ year timescales, and given that the scaling of universe growth is exponential with time, the number of universes with regular-brain observers will be exponentially larger than Boltzmann brains. This applies to both CNS1 and the revised version we have presented, the latter with an even larger profusion of BHs made in a relatively short $10^{14}$ yr period. Note that eternal inflation, which is invoked as an alternative to CNS (Sec. III.1), generally leads to the Boltzmann brain problem Albrecht and Sorbo (2004); Bousso and Freivogel (2007); Linde (2007); Carroll (2017).
@@ Acknowledgments
We thank Lee Smolin for conversations and comments on this work and thank Robert Brandenberger and Aaron Vincent for discussions. This work was supported by the Arthur B. McDonald Canadian Astroparticle Physics Research Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Canada Foundation for Innovation. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario.
Appendix A Critical commentary on tentative predictions of cosmological natural selection
In the original version of CNS, Smolin proposed three tests that were suggested as ways to falsify the hypothesis. If these tests are taken at face value it could be argued that the hypothesis has been falsified. In Section V we have already discussed why, in our opinion, the predictions laid out in CNS1 are not necessarily ideal tests. Moreover, we think our findings in this work have provided some new ways to test the theory. Our major objection to the original tests is that the global fitness of the universe need not increase by naively varying a single parameter to increase BH production, which is in agreement with results from evolutionary biology Stearns (2000). Nevertheless, we think it is instructive to review the predictions laid out in Ref. Smolin (1992), and comment on their status vis-a-vis current astrophysical and cosmological data. The CNS1 predictions advocated in Ref. Smolin (1994, 1997) are as follows:
(i) Neutron stars must be composed of kaon-condensates, which soften the high-density equation of state and hence lower the maximum mass of stable NSs, $M^{\rm NS}_{\rm max}$. The smaller $M^{\rm NS}_{\rm max}$ is, the wider the range of progenitor masses that end up as BHs as opposed to NSs. The strange quark mass is then optimal in our universe [^3] Apart from this prediction, CNS may also explain the existence of the heavier fermion generations as relics of early (literal) generations of self-reproducing universes mutating toward a set of fermion masses optimal for production of stars and BHs Nambu (1985); Smolin (1992). as variations in it could vary $M^{\rm NS}_{\rm max}$ without influencing massive star formation and supernovae. In Ref. Smolin (2006) $M^{\rm NS}_{\rm max}$ is predicted as about 1.6 $M_{\odot}$.
@@ Appendix A Critical commentary on tentative predictions of cosmological natural selection
(iii) Early star formation must be suppressed. Organic chemistry is required for the cooling of star-forming molecular hydrogen gas clouds from which BH progenitors arise. In particular, radiation from molecular vibrations of CO dominantly cool the gas clouds, and carbon dust and ice shield the clouds from ultraviolet light emitted by massive stars. If an alternative channel existed, it would would have operated at high redshifts when carbon and oxygen abundances were much smaller, and produced many more core-collapse supernovae than observed. Thus the neutron-proton mass difference, being smaller than nuclear binding energies in our universe, is selected to pave the way for star formation but only at $t>10^{6}$ yr.
Now we provide our remarks. Prediction (i), on the face of it, is in tension with the observation of NSs heavier than 2 $M_{\odot}$; the heaviest at the time of writing is estimated to weigh 2.35 $\pm$ 0.17 $M_{\odot}$ Romani et al. (2022). This prediction/bound for heavy NSs was in any case blurry. The extra span in mass range of black holes made by core-collapse of main-sequence stars to kaon-condensate (or hyperon) stars would arguably result in a small gain in black holes formed, compared to the total range of astrophysical black hole masses, $\mathcal{O}(10^{0}-10^{2})~{}M_{\odot}$, which is further impacted by the effects of stellar metallicity Spera et al. (2015); Sicilia et al. (2022). Moreover, it is not clear that nature’s actual dynamics for forming BHs and NSs in core collapse supernovae would result in more BHs being formed if the maximum NS mass is lowered. There is a deficit (“mass gap”) observed in 2--5 solar mass compact objects in x-ray binary systems Farr et al. (2011), which is in tension with models of core collapse supernovae that predict a continuous population of NSs/BHs produced in this mass range. If core collapse supernovae, owing to the still-emerging details of their dynamical collapse, do not form compact objects in this mass range, then the maximum mass of NSs – itself set by the unknown high-density equation of state of nuclear matter – may be irrelevant for determining the number of black holes produced.
@@
Finally, we discuss one notable objection to the paradigm of CNS, which was raised by Vilenkin Vilenkin (2006). In the far future, when the universe may be described by de Sitter space, semiclassical descriptions of gravity with instantons imply that BHs will be nucleated via quantum tunneling. The nucleation rate goes as $\exp[-M_{\rm BH}/T]$, with $T=H/2\pi=\sqrt{\Lambda_{\rm cc}/3}/2\pi$ the Gibbons-Hawking temperature of the quantum fields in de Sitter space, where $H$ is the Hubble rate and $\Lambda_{\rm cc}$ is the CC. Thus by increasing $\Lambda_{\rm cc}$ to more than the observed value, the number of BHs formed is exponentially increased, apparently falsifying CNS. Smolin’s response Smolin (2006) includes the arguments that currently known physics cannot be extrapolated to much vaster scales in distance, time, and energy, that CNS concerns itself with explaining the universe as we observe it today, and that random fluctuations can always produce entities that outnumber their equivalents formed through an ordering principle, e.g., DNA sequences formed outside reproducing organisms and Poincare recurrence-induced Boltzmann brains, due to which care must be taken in choosing the ensemble in which we want to show that an entity is “typical”. (These arguments also apply to the criticism of CNS that blue-tilting the spectral index of primordial fluctuations would enhance BH production Silk (1997).)

### v2

@@ II Summary
    Here we will propose that most of our universe’s CNS-associated black hole production may actually occur through dark matter converting old helium white dwarfs (WDs) into about $0.1~{}{M_{\odot}}$ black holes. This can transmute nearly all baryons in galaxies into black holes over 100 trillion year timescales.
@@
Figure 1: The total entropy production in our universe with associated uncertainties (solid blue lines) as a function of time, using the historic star formation rate and uncertainties reported in Ref. Hopkins and Beacom (2006), and extrapolated to future epochs using future star formation Adams and Laughlin (1997) as detailed in Sec. IV.2. The dashed black curve shows the amount of galactic stellar material that would be turned into black holes for dark matter with the mass and per-nucleon cross section indicated. This result is shown both in terms of the number fraction (labelled “num.”) and mass fraction (labelled “mass”) of stars converted. In the absence of dark matter conversion (“no DM conv.”), baryons predominantly form black holes through core collapse supernovae in a roughly constant proportion Sicilia et al. (2022). The calculations in Secs. IV.1-IV.2 account for dark matter conversion to black holes of $0.1-0.5~{}M_{\odot}$ He white dwarfs, which are the predominant site of baryons in galaxies after star formation ceases. Dark matter-induced collapse will also apply to higher mass white dwarfs, neutron stars Gould et al. (1990); Kouvaris (2008); Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010); McDermott et al. (2012); Kouvaris and Tinyakov (2011); Kouvaris (2012); Güver et al. (2014); Bertoni et al. (2013); Bramante et al. (2013); Bell et al. (2013); Bramante and Linden (2014); Bramante (2015); Bramante and Linden (2016); Bramante et al. (2018); Garani et al. (2019); Kouvaris et al. (2018); Kopp et al. (2018); Acevedo and Bramante (2019); East and Lehner (2019); Tsai et al. (2021); Dasgupta et al. (2020); Garani et al. (2021, 2022); Giffin et al. (2022); Ray (2023), planets Acevedo et al. (2021), and other celestial objects, a sub-dominant contribution not included here.
@@ III.1 Cosmological natural selection
Cosmological natural selection is based on the notion that typical universes are those that reproduce best. As detailed in Ref. Smolin (1992), the reproduction of universes is hypothesized to occur via black hole formation, with the assumption that a causally disconnected universe is born beyond the horizon from a bouncing singularity. (This idea is different from “the universe as a black hole” Pathria (1972) where an equivalence is suggested between the FRW metric of an internal observer and the Schwarzschwild metric of an external one.) In each new universe, physical constants slightly vary, presumably due to a dense discretuum of vacua whose minima vary by a small amount due to the bounce dynamics (such as might be seen in something like a string theory landscape). This results in a sum of universes whose greater part is optimized to create black holes. Ref. Smolin (1992) points out, for instance, that if the neutron-proton mass difference were negative, radiative cooling of neutron (as opposed to hydrogen) gas clouds would be too slow and suppress star formation; if it were higher than the observed 1.29 MeV, nucleosynthesis of elements heavier than hydrogen would be slower, which in turn would subdue the formation of later-generation stars from recycled supernova-ejected stellar material. If it were too high, stable deuterium, and thence heavier nuclides, may never be formed in Big Bang nucleosynthesis. Analogous to the CNS mechanism is Darwinian natural selection, with the slight changes in genetic code paralleling the slight changes in physical constants.
@@ III.4 Linking entropy and black hole production
We now propose the following update to CNS in Smolin’s original form, which for clarity here we will call CNS1. In CNS1, about 1% of baryonic mass ends up inside black holes (through conventional core-collapse of superheavy stars); in the new version, more than 75% of the galactic baryonic mass becomes black holes. [^1] As discussed later, this number is likely higher once black hole conversion of $>0.5$ solar mass stars and NSs are accounted for. In CNS1 the criteria by which the framework is evaluated is tied to known black hole production mechanisms, but does not tie into the dynamic timescale and production of entropy in our universe; in the CNS proposed here, timescales for entropy production and creation of new universes come out to be about the same, and this is quantified using stelliferous entropy. Finally, in our proposed reframing of CNS, dark matter plays a central and intricate role in converting stars into black holes.
@@ IV Conversion of white dwarfs to black holes via dark matter
In this section, we estimate the timescale over which low-mass helium WDs, end products of low-mass main sequence stars, would be converted into black holes by heavy asymmetric dark matter. We assume that dark matter is a Dirac fermion, although as we will see shortly our treatment would apply to bosonic dark matter states as well. For further aspects of dark matter capture in compact stars see the review in Ref. Bramante and Raj (2024). The mass of Dirac fermionic dark matter most relevant to this study will be $m_{\chi}\gtrsim 10^{6}$ GeV. The cosmological production of such heavy asymmetric dark matter has been detailed in a number of prior works such as Refs. Benakli et al. (1999); Chung et al. (2001); Bramante and Unwin (2017); Baker et al. (2020); Kramer et al. (2021); Asadi et al. (2021), and has been associated with, e.g., moduli decay, other sources of entropy-production or early matter domination, first-order phase-transitions, and non-equilibrium production.
@@ IV.1 Dark matter-white dwarf dynamics
where this expression sums over rates for dark matter to be captured after $n$ scatters, given by Bramante et al. (2017)
@@
We now comment on the timescales for dark matter to accumulate in and convert compact stars into black holes. First, we note that the above expression for the evaporation timescale should break down at sufficiently small black hole mass, since this introduces a high-energy regime of black hole evaporation which is still being studied Page (2005). Secondly, it will be useful for future numerical simulations of dark matter in compact stars to validate or improve on the above timescale estimates. One such study has been conducted for a small black hole formed from dark matter inside a neutron star, which found reasonable agreement with the late-stage millisecond-long Bondi accumulation time predicted by the above formulae East and Lehner (2019).
@@
IV.2 Entropy production and black hole formation
To estimate the integrated entropy production rate shown in Figure 1 at time $t$ we follow Ref. Bousso et al. (2007), with some modifications. We use the Kroupa IMF Kroupa (2002) given in Sec. III.3, $\xi_{\rm IMF}(M_{\star})$, which weights the entropy production rate per star, $d^{2}s/dN_{\star}dt=L_{\star}/T_{\rm eff}$, where the main sequence star luminosity $L_{\star}\propto M_{\star}^{3.5}$ and the temperature of reprocessed starlight $T_{\rm eff}=20$ meV. This is then convolved with the redshift-dependent star formation rate, $\dot{\rho}_{\star}$, and integrated over the birth times $t^{\prime}$ of stars born prior to $t$ and volume $V_{c}$:
@@ V.1 Alternative CNS measures and dynamics
This brings us to the question of whether non-stellar sources of entropy increase might be considered. In our work we have used the entropy sourced by reprocessed starlight as the predominant non-horizon entropy source. But far more entropy than this is sourced by horizons, $i.e.$ of both black holes and the cosmic horizon of the observable universe. As in Ref. Bousso et al. (2007) we have not considered this horizon entropy in the main text, as its relevance to observers and the birth of universes is unclear. If horizon entropy is considered for some reason, our universe as we know it is enough to convert most of the mass of entropy sources to new universes. This is because in $10^{33}$ yr most baryonic mass would have gone into making galactic mass black holes through dynamical accumulation onto the supermassive black holes (SMBHs) at the centers of galaxies Adams and Laughlin (1997). These eventual SMBHs would be the major sources of horizon entropy. In that scenario, star formation plays a key role, but the role of dark matter is unclear. But if it is part of the equation, models of dark matter seeding SMBHs Feng et al. (2021) become significant. The black hole count may also be maximized in these scenarios. In addition, one could have super-massive PBHs that grow faster than regular SMBHs and contributing to the majority of entropy, although these could only make up a small fraction of dark matter owing to constraints from their accretion Green and Kavanagh (2021). Also, these dark matter-as-black holes versions of CNS do not explain star formation.
@@ V.2 Perspective on reccurrence
To wrap up, we discuss measure problems and their relation to CNS. As a prior example of measure problems, let us consider Weinberg’s work on understanding the cosmological constant. In formulating his anthropic bound on the CC, Weinberg had assumed that the fraction of baryons that ends up in galaxies is another parameter (in addition to the CC) that scans on some landscape of vacua, as the anthropically conditioned probability distribution is proportional to the number of observers. But he had raised the possibility that the number of baryons may itself scan Weinberg (2005), in which case one cannot know that our universe is typical. Analogously, one parameter that could scan in the CNS framework is the fraction of baryonic mass that converts to black holes. As we have noted at the outset of our discussion, scanning parameters in this way does not necessarily lead to increased universe fitness.
However, in the context of cosmologies whose future looks similar to their past, let us note that CNS appears to generically circumvent the Boltzmann brain problem. Over timescales of $\gg 10^{10^{66}}$ yr, thermodynamic fluctuations could create low-entropy states that mimic conscious observers with false memories of a briefer evolution of the universe Carroll (2017). The problem is that these “Boltzmann brains” would far outnumber “regular brains” formed through better known evolutionary pathways, implying we must ask whether we are the former. The solution would be to construct a cosmology where Boltzmann brains are atypical. In CNS as we have outlined, the formation of new universes occurs over $10^{14}$ year timescales, and given that the scaling of universe growth is exponential with time, the number of universes with regular-brain observers will be exponentially larger than Boltzmann brains. This applies to both CNS1 and the revised version we have presented, the latter with an even larger profusion of black holes made in a relatively short $10^{14}$ yr period. Note that eternal inflation, which is invoked as an alternative to CNS (Sec. III.1), generally leads to the Boltzmann brain problem Albrecht and Sorbo (2004); Bousso and Freivogel (2007); Linde (2007); Carroll (2017).
@@ Acknowledgments
We thank Lee Smolin for conversations and comments on this work, and Robert Brandenberger, Shantanu Desai and Aaron Vincent for discussions. This work was supported by the Arthur B. McDonald Canadian Astroparticle Physics Research Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Canada Foundation for Innovation. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario.
Appendix A Critical commentary on cosmological natural selection
In the original version of CNS, Smolin proposed three tests that were suggested as ways to falsify the hypothesis. If these tests are taken at face value it could be argued that the hypothesis has been falsified. In Section V we have already discussed why, in our opinion, the predictions laid out in CNS1 are not necessarily ideal tests. Moreover, we think our findings in this work have provided some new ways to test the theory. Our major objection to the original tests is that the global fitness of the universe need not increase by naively varying a single parameter to increase black hole production, which is in agreement with results from evolutionary biology Stearns (2000). Nevertheless, we think it is instructive to review the predictions laid out in Ref. Smolin (1992), and comment on their status vis-a-vis current astrophysical and cosmological data. The CNS1 predictions advocated in Ref. Smolin (1994, 1997) are as follows:
(i) Neutron stars must be composed of kaon-condensates, which soften the high-density equation of state and hence lower the maximum mass of stable NSs, $M^{\rm NS}_{\rm max}$. The smaller $M^{\rm NS}_{\rm max}$ is, the wider the range of progenitor masses that end up as black holes as opposed to NSs. The strange quark mass is then optimal in our universe [^3] Apart from this prediction, CNS may also explain the existence of the heavier fermion generations as relics of early (literal) generations of self-reproducing universes mutating toward a set of fermion masses optimal for production of stars and BHs Nambu (1985); Smolin (1992). as variations in it could vary $M^{\rm NS}_{\rm max}$ without influencing massive star formation and supernovae. In Ref. Smolin (2006) $M^{\rm NS}_{\rm max}$ is predicted as about 1.6 $M_{\odot}$.
@@ Appendix A Critical commentary on cosmological natural selection
(iii) Early star formation must be suppressed. Organic chemistry is required for the cooling of star-forming molecular hydrogen gas clouds from which black hole progenitors arise. In particular, radiation from molecular vibrations of CO dominantly cool the gas clouds, and carbon dust and ice shield the clouds from ultraviolet light emitted by massive stars. If an alternative channel existed, it would would have operated at high redshifts when carbon and oxygen abundances were much smaller, and produced many more core-collapse supernovae than observed. Thus the neutron-proton mass difference, being smaller than nuclear binding energies in our universe, is selected to pave the way for star formation but only at $t>10^{6}$ yr.
Now we provide our remarks. Prediction (i), on the face of it, is in tension with the observation of NSs heavier than 2 $M_{\odot}$; the heaviest at the time of writing is estimated to weigh 2.35 $\pm$ 0.17 $M_{\odot}$ Romani et al. (2022). This prediction/bound for heavy NSs was in any case blurry. The extra span in mass range of black holes made by core-collapse of main-sequence stars to kaon-condensate (or hyperon) stars would arguably result in a small gain in black holes formed, compared to the total range of astrophysical black hole masses, $\mathcal{O}(10^{0}-10^{2})~{}M_{\odot}$, which is further impacted by the effects of stellar metallicity Spera et al. (2015); Sicilia et al. (2022). Moreover, it is not clear that nature’s actual dynamics for forming black holes and NSs in core collapse supernovae would result in more black holes being formed if the maximum NS mass is lowered. There is a deficit (“mass gap”) observed in 2--5 solar mass compact objects in x-ray binary systems Farr et al. (2011), which is in tension with models of core collapse supernovae that predict a continuous population of NSs/black holes produced in this mass range. If core collapse supernovae, owing to the still-emerging details of their dynamical collapse, do not form compact objects in this mass range, then the maximum mass of NSs – itself set by the unknown high-density equation of state of nuclear matter – may be irrelevant for determining the number of black holes produced.
@@
Finally, we discuss one notable objection to the paradigm of CNS, which was raised by Vilenkin Vilenkin (2006). In the far future, when the universe may be described by de Sitter space, semiclassical descriptions of gravity with instantons imply that black holes will be nucleated via quantum tunneling. The nucleation rate goes as $\exp[-M_{\rm BH}/T]$, with $T=H/2\pi=\sqrt{\Lambda_{\rm cc}/3}/2\pi$ the Gibbons-Hawking temperature of the quantum fields in de Sitter space, where $H$ is the Hubble rate and $\Lambda_{\rm cc}$ is the CC. Thus by increasing $\Lambda_{\rm cc}$ to more than the observed value, the number of black holes formed is exponentially increased, apparently falsifying CNS. Smolin’s response Smolin (2006) includes the arguments that currently known physics cannot be extrapolated to much vaster scales in distance, time, and energy, that CNS concerns itself with explaining the universe as we observe it today, and that random fluctuations can always produce entities that outnumber their equivalents formed through an ordering principle, e.g., DNA sequences formed outside reproducing organisms and Poincare recurrence-induced Boltzmann brains, due to which care must be taken in choosing the ensemble in which we want to show that an entity is “typical”. (These arguments also apply to the criticism of CNS that blue-tilting the spectral index of primordial fluctuations would enhance black hole production Silk (1997).)